AN OPTICAL-OPTICAL DOUBLE-RESONANCE SPECTROSCOPIC STUDY OF FOUR ION-PAIR STATES OF CLF AND IDENTIFICATION OF THE CLF(A) VALENCE STATE

\underline{V. A. ALEKSEEV}^{a} and D. W. SETSER, Chemistry Department, Kansas State University, Manhattan, KS 66506.

Four ion-pair states of ClF, E 0^{+}(^{3}P_{2}), f 0^{+}(^{3}P_{0}), β 1(^{3}P_{2}) and G 1(^{3}P_{1}), have been identified by sequental excitation via the B ^{3}Π_{0+} valence state. Excitation spectra and fluorescence spectra have been acquired, which permit the assignment of vibrational energies and rotational constants to 31 levels. These four ion-pair states exhibit complex homogeneous and heterogeneous interactions and neither the vibrational nor rotational constants are regular with increasing vibrational quantum number. The vibrational and rotational constants of the A ^{3}Π_{1} state were identified from the low resolution β − A emission spectra and the dissociation limits of ClF (B ^{3}Π_{0+}) and ClF(A ^{3}Π_{1}) are discussed. The dissociation energy of ClF (X) is confirmed to be 21110 cm^{−1}. Some qualitative information also was obtained about the D′ 2(^{3}P_{2}) and A′ ^{3}Π_{2} states.

aPermanent address: Photonics Department, Institute of Physics, St.Petersburg University, 198904 St. Petersburg, RUSSIA